# Searching for coupled, hyperlight scalars across cosmic history

### v1

@@ IV.2.1 Effects of the matter potential
In the late Universe, matter-induced dynamics are dictated by the in-medium potential Eq. 38 and parameterized by the dimensionless combination $\mathcal{D}$ from Eq. 50. Section A.1 shows that the field begins evolving before matter-radiation equality if $|\mathcal{D}|\gtrsim 1$—that is, if the in-medium mass is comparable in size to $H$. Because the CMB visibility function peaks around $a_{\star}/a_{\mathrm{eq}}\approx 3$ (where $a_{\mathrm{eq}}$ is the scale factor of matter-radiation equality), the fundamental constants would evolve nonnegligibly during recombination for such values of the couplings. These dynamics would be inconsistent with the treatment of Ref. [60] (and prior studies of the CMB with varying constants); they are not precluded a priori but are simply beyond the scope of the present work. In the regime $|\mathcal{D}|\lesssim 1$, the approximate solution Eq. 51 sets an estimated shift in fundamental constants $\lambda$ between early times (post annihilation) and recombination of
$\left|\lambda_{\star}/\Delta\lambda_{i}-1\right|=\left|\bar{\varphi}(a_{\star})^{2}/\bar{\varphi}_{i}^{2}-1\right|$ $\approx\left|(Q_{b})_{\lambda}d_{\lambda}^{(2)}\right|/3.$ (95)
Requiring the early-time shift in $\alpha$ or $m_{e}$ to remain constant at the $10\%$ level then requires $|d_{\lambda}^{(2)}|\lesssim 500$.
As discussed in Section II.5, matter-induced dynamics can also modify the relationship between the scalar’s early-time value $\bar{\varphi}_{i}$ and its present-day abundance $F_{\varphi}$, which we take to be given exactly by the free-evolution result, Eq. 53, in this section. The total deviation is the sum of that accumulated through Eq. 51 evaluated at $a(t_{\mathrm{osc}})$ and the small deviation from an $a^{-3/2}$ oscillation envelope at $t>t_{\mathrm{osc}}$ (see Section A.1). The former effect is more pronounced for lighter fields that begin oscillating later. Requiring the matter-free result Eq. 53 to be accurate at the $10\%$ level constricts the regime of validity by a further factor of two to five for scalars with masses $10^{-29}\leavevmode\nobreak\ \mathrm{eV}$ to $10^{-31}\leavevmode\nobreak\ \mathrm{eV}$.
@@ IV.2.2 Cosmological constraints
Figure 4: Posterior distributions over a hyperlight scalar field’s present energy density relative to that of cold dark matter, $F_{\varphi}=\omega_{\varphi}/\omega_{c}$, and its quadratic couplings to the electron (left) and the photon (right). The diagonal panels depict one-dimensional marginal distributions, and the lower left panel depicts $1$ and $2\sigma$ mass levels of the joint, two-dimensional posterior from parameter inference using Planck 2018 data alone (blue) and combined with BAO data (red). Grey lines depict $2\sigma$ limits from quasar absorption lines (derived in Section III.3) for a set of scalar masses $m_{\varphi}$ as labeled. For the early-time constraints, the scalar mass is fixed to $m_{\varphi}=10^{-30}\leavevmode\nobreak\ \mathrm{eV}$, but these results do not change qualitatively within the range $10^{-32}\lesssim m_{\varphi}/\mathrm{eV}\lesssim 10^{-29}$. Note that parameter inference for the early-time constraints place flat priors on $F_{\varphi}$ and the early-time parameter value $\lambda_{i}$; the increase in posterior density with $F_{\varphi}$ at low values is the result of plotting the distribution over $\log_{10}F_{\varphi}$.
@@
Because Fig. 4 takes uniform priors over $F_{\varphi}$, the marginalized posteriors over $\log_{10}F_{\varphi}$ grow exponentially at small $F_{\varphi}$; they are then sharply cut off at values near half a percent, a feature shared with results for scalars with purely gravitational interactions [87] that derives solely from the disfavored effects of hyperlight scalar fields on structure growth. Namely, the small-$F_{\varphi}$ tail of the marginalized posterior is driven mostly by the prior, while the sharp decline of posterior probability beyond $F_{\varphi}=\mathcal{O}(10^{-2})$ is driven by the data. Further, because uniform priors disfavor extremely small values for $F_{\varphi}$, the bulk of each posterior always lies within the regime of couplings where matter effects on the scalar’s dynamics (Section IV.2.1) are negligible. This holds even for the broad, Planck-only constraints on the electron coupling, for which $95\%$ of posterior samples lie at values of $d_{m_{e}}^{(2)}$ above $-500$. We discuss the impact of alternative priors in Section B.1.
@@ IV.3 Joint constraints
Figure 5: Exclusion regions for the quadratic couplings of a new scalar to the electron (left column) and photon (right column) that are positive (top row) or negative (bottom row). Results are shown for a fiducial abundance of the scalar (relative to that of CDM) of $F_{\varphi}\equiv\bar{\rho}_{\varphi,0}/\bar{\rho}_{c,0}=10^{-2}$, which is roughly the largest allowed in our analysis for the mass range depicted here [60]. The bounds on quadratic couplings, however, are mostly driven by the effect on the early-time values of $\alpha$ and $m_{e}$, as are the constraints from quasars (Section III.3) and BBN (Section IV.1). The CMB data are from the 2018 Planck release, which are presented in combination with BAO data from SDSS [175, 176] and 6dFGS [177] (dark blue), BAO data from DESI [172] (light blue), and SNe data from either DES [174] (dashed) or Pantheon+ [178, 179] (dotted). Observe that the combination with DESI yields an apparent preference for nonzero, positive electron-mass couplings at $2\sigma$ that is not particularly consistent with the other dataset combinations, all of which are consistent with zero; see Section IV.2 and Ref. [60] for exposition on the cosmological sensitivity of these measurements. At masses higher than those depicted, the scalar deviates from its initial condition by more than $10\%$ at recombination, a regime which requires a dedicated analysis. Results approximately account for the impact on the relationship between mass and initial misalignment for scalars that begin oscillating around or before matter-radiation equality. Results are plotted on a symmetric logarithmic scale that interpolates linearly about zero between values of ±$\pm 10^{-2}$ via a transformation using an inverse hyperbolic function ($\mathrm{arcsinh}$).
@@ V.2 Local and late-time searches
In Section III we extend the analysis of these late-time probes to handle several qualitatively distinct features of hyperlight scalars, with results summarized in Fig. 2. Measurements long after a hyperlight scalar begins oscillating are plagued by its extremely long oscillation period: the statistical possibility that the measurement occurs near a node of the signal steeply penalizes the reach of single, finite-duration probes. As a result, atomic clocks and tests of the equivalence principle are limited in constraining power. Section III.1 argues that the incurred penalty depends on the signal’s scaling with phase about its nodes. Equivalence principle tests (Section III.4) for quadratic couplings are therefore penalized by two powers of mass times inverse experimental duration, while atomic clock and pulsar timing arrays (Section III.6), which essentially probe the drift rate of fundamental constants, incur a penalty of one power for both linear and quadratic couplings. Combinations of measurements can mitigate this penalty only if spaced in time by a substantial fraction of the scalar’s oscillation period because of the greatly suppressed chance that the scalar’s amplitudes sampled at random points in its oscillation all occur at nodes. Constraints on varying constants from quasar absorption lines (Section III.3) span redshifts as high as $7$ and therefore remain competitive when analyzed as an ensemble (Fig. 3), even if during epochs when the field oscillates, setting the leading constraints at the lowest masses.
@@ V.4 Future directions
To sum up, the phenomenology of scalar fields coupled to the SM is rich, and cosmological observations offer a powerful probe of unique regimes of parameter space, both via precise sensitivity to altered microphysics in the early Universe and through the gravitational impact of the new fields. Moreover, future opportunities to search for coupled scalars (hyperlight or otherwise) are myriad [189], including nuclear clocks [190, 191, 192, 193], high-resolution CMB observations [194, 195], the Lyman-$\alpha$ forest [196], and absorption spectra measurements from quasars [197] and emission line galaxies [198]. In the near term, further data from DESI and other future galaxy surveys [199, 200, 201, 202, 203] and supernova surveys [204, 205, 206] should improve limits on electron couplings and clarify the discordant results from current datasets. Studying the interesting dynamical regimes of strong couplings and heavier masses, as well as couplings to the strong sector, would paint a complete picture of new hyperlight, interacting scalars, which may well be first revealed through their cosmological signatures in the wealth of upcoming data.
@@ A.1 After annihilation
Figure 6: Solutions to Eq. 96 for a quadratically coupled scalar with a nonrelativistic matter potential, with values for $(Q_{b})_{\lambda}d_{\lambda}^{(2)}$ varying by color. Vertical black lines indicate the moment of recombination (i.e., for the fiducial case that the fundamental constants do not change with time).
@@
We now determine what maximum $|d_{\lambda}^{(2)}|$ permits the scalar to be sufficiently stable that we may treat the variations in fundamental constants as themselves time independent before recombination. For $|(Q_{b})_{\lambda}d_{\lambda}^{(2)}|\lesssim 1$, most of the scalar’s evolution takes place in the short period between matter-radiation equality and recombination and can be understood with the matter-era solution; we empirically find $\varphi/\varphi_{i}\approx(1+2a/3a_{\mathrm{eq}})^{\beta_{+}}$ provides an excellent approximation, yielding Eq. 51 at small $\mathcal{D}$. For Planck’s preferred $\Lambda$CDM parameters [99], the scale factor of recombination relative to that at equality is $a_{\star}/a_{\mathrm{eq}}\approx 3$ and $\omega_{c}/\omega_{b}\approx 5.4$. In the small-$\mathcal{D}$ limit, $\beta_{+}\approx-2\mathcal{D}/3$ and the shift in $\lambda$ varies by recombination by an amount given by Eq. 95.
@@
where $F_{\varphi}$ is the scalar-CDM ratio given by Eq. 53, i.e., for an uncoupled scalar. Equation 102 encodes the correction to the relationship between the early-time parameter shift $\Delta\lambda_{i}/\lambda(0)$, $d_{\lambda}^{(2)}$, and the late-time abundance $F_{\varphi}$ for an uncoupled scalar, Eq. 58. As explained in Section IV.3, because these effects are relevant only at large $d_{\lambda}^{(2)}$ but cosmological data place relatively strong constraints on $\Delta\lambda_{i}/\lambda(0)$, they are only relevant in the regime in which the scalar’s abundance is negligibly small. Any changes to the scalar’s dynamics are then themselves negligible. Equation 102 still encodes the correction between $F_{\varphi}$ (taken to be defined as the combination in Eq. 58) and the present-day scalar abundance relevant to, e.g., laboratory probes. But these effects are again small in the parameter space we consider.
@@ A.2 Before annihilation
Figure 7: Solutions to Eq. 103 for a scalar coupled quadratically to the electron (left) and photon (right), with couplings $d_{\lambda}^{(2)}$ varying by color.
@@
Oscillatory solutions for positive couplings again switch to exponentially growing ones for negative couplings. For large negative couplings $-d_{m_{e}}^{(2)}\gtrsim 100$, numerical solutions determine that the scalar refreezes at a value $\sim e^{3\sqrt{|\mathcal{D}_{m_{e}}|}}$; the scaling with $\sqrt{|\mathcal{D}_{m_{e}}|}$ is as expected from Eq. 104, while the numerical coefficient reflects the particular dynamics of refreezing as the in-medium mass drops exponentially during annihilation. At the classical level, there is no theoretical issue with the scalar having an arbitrarily small value before reheating that is substantially amplified to its post-annihilation value. However, during inflation, a quantized scalar scalar field with $m_{\varphi}<H_{\text{inf}}$ acquires an average energy density of order $\sim H_{\mathrm{inf}}^{2}/m_{\varphi}>H_{\text{inf}}$ [218]. Heuristically, this is because an inflationary Universe has a IR cutoff at the inflationary scale $H_{\text{inf}}$, meaning that the smallest physically meaningful scale is $\mathcal{O}(H_{\text{inf}})$. At the very minimum, inflation had to occur in time for reheating to complete at a temperature $T_{\mathrm{RH}}\approx 4\leavevmode\nobreak\ \mathrm{MeV}$ [75, 76, 77, 78, 79, 80, 81], limiting $H_{\mathrm{inf}}\gtrsim 10^{-14}\leavevmode\nobreak\ \mathrm{eV}$. Requiring that $|\varphi_{\text{RH}}|>H_{\text{inf}}/\sqrt{2}M_{\mathrm{pl}}$ implies
@@
The scalar’s effective mass induced by the coupling to photons is proportional to $H$ at all times after reheating.999 Technically, the rest of the charged SM content would contribute to the interacting free energy at higher temperatures. The SM presumably also generates an effective potential for the scalar before it fully thermalizes, but it would not be captured by the thermal field theory calculations Section II.3.2 relies on. We set aside these complex and uncertain details for the sake of simplicity. Now defining $\mathcal{D}_{e}=25\alpha(0)d_{e}^{(2)}/8\pi g_{\star}$, the solution under the photon coupling is
@@
where $\Delta\alpha_{i}$ is the change in $\alpha$ post-annihilation that is the initial condition for the dynamics leading up to recombination. Again taking the minimum reheat temperature $T_{\mathrm{RH}}\approx 4\leavevmode\nobreak\ \mathrm{MeV}$, this bound amounts to $d_{e}^{(2)}\lesssim 360$; for instantaneous reheating after high-scale inflation, the bound tightens by a factor $\sim 20$.
@@ B.1 Impact of priors on marginalized constraints
While any choice of prior on phenomenological parameters ultimately lacks direct theoretical motivation, even that for fundamental parameters is ambiguous. For instance, from the perspective of effective field theory one might consider the dimensionless parameter $d_{\lambda}^{(2)}$ an order-unity Wilson coefficient suitable for a uniform prior. On the other hand, $M_{\mathrm{pl}}/d_{\lambda}^{(2)}$ may be better described as a new-physics scale for a UV completion of the model, and one might choose for it a log-uniform prior in order to avoid arbitrarily selecting a preferred energy scale. [That is, a uniform prior between $a$ and $b$ identifies $(b-a)/2$ as a “special” value, whereas a log-uniform prior allocates equal weight to each decade within the allowed limits.] A choice for $\bar{\varphi}_{i}$ could also be informed by the choice for $d_{\lambda}^{(2)}$—perhaps $\bar{\varphi}_{i}$ is itself a characteristic energy scale in the theory, or derives from some other physics that generates the initial conditions, or so on. Posteriors, of course, are most useful when they are least sensitive to prior choices.
@@
Compared to a uniform prior over $F_{\varphi}$ and $\lambda_{i}/\lambda_{0}$, this prior penalizes large $F_{\varphi}$ quadratically. If the posteriors over $d_{\lambda}^{(2)}$ are largely driven by the likelihood’s dependence on $\lambda_{i}/\lambda_{0}$, we would expect such a prior to drive the marginal posteriors over the quadratic couplings $d_{\lambda}^{(2)}$ to values even larger in magnitude. Priors must be proper, and the chosen limits on, e.g., $\bar{\varphi}_{i}$ determine the relative prior weight allocated percent-level values of $F_{\varphi}$—i.e., those that were allowed in the results taking uniform phenomenological priors. Specifically, we impose the effective prior Eq. 113 with boundaries corresponding to $d_{\lambda}^{(2)}\sim\mathcal{U}(-10^{3},10^{3})$ and $\log_{10}\bar{\varphi}_{i}\sim\mathcal{U}(-2,0)$, where $\mathcal{U}$ denotes a uniform distribution. We additionally restrict $10^{-4}\leq F_{\varphi}\leq 0.3$ in order to roughly match the range of values covered by the posteriors using uniform, phenomenological priors.

### v2

@@ IV.2.1 Effects of the matter potential
In the late Universe, matter-induced dynamics are dictated by the in-medium potential Eq. 38 and parametrized by the dimensionless combination $\mathcal{D}$ from Eq. 50. Section A.1 shows that the field begins evolving before matter-radiation equality if $|\mathcal{D}|\gtrsim 1$—that is, if the in-medium mass is comparable in size to $H$. Because the CMB visibility function peaks around $a_{\star}/a_{\mathrm{eq}}\approx 3$ (where $a_{\mathrm{eq}}$ is the scale factor of matter-radiation equality), the fundamental constants would evolve nonnegligibly during recombination for such values of the couplings. These dynamics would be inconsistent with the treatment of Ref. [60] (and prior studies of the CMB with varying constants); they are not precluded a priori but are simply beyond the scope of the present work. In the regime $|\mathcal{D}|\lesssim 1$, the approximate solution Eq. 51 sets an estimated shift in fundamental constants $\lambda$ between early times (post annihilation) and recombination of
$|\displaystyle\left|\lambda_{\star}/\Delta\lambda_{i}-1\right|=\left|\bar{\varphi}(a_{\star})^{2}/\bar{\varphi}_{i}^{2}-1\right|$$\approx\left|(Q_{b})_{\lambda}d_{\lambda}^{(2)}\right|/3.$ (95)
[…]
@@ IV.2.2 Cosmological constraints
Figure 4: Posterior distributions over a hyperlight scalar field’s present energy density relative to that of cold dark matter, $F_{\varphi}=\omega_{\varphi}/\omega_{c}$, and its quadratic couplings to the electron (left) and the photon (right). The diagonal panels depict one-dimensional marginal distributions, and the lower left panel depicts 111 and $2\sigma$ mass levels of the joint, two-dimensional posterior from parameter inference using Planck 2018 data alone (blue) and combined with BAO data (red). Grey lines depict $2\sigma$ limits from quasar absorption lines (derived in Section III.3) for a set of scalar masses $m_{\varphi}$ as labeled. For the early-time constraints, the scalar mass is fixed to $m_{\varphi}=10^{-30}\leavevmode\nobreak\ \mathrm{eV}$, but these results do not change qualitatively within the range $10^{-32}\lesssim m_{\varphi}/\mathrm{eV}\lesssim 10^{-29}$. Note that parameter inference for the early-time constraints place flat priors on $F_{\varphi}$ and the early-time parameter value $\lambda_{i}$; the increase in posterior density with $F_{\varphi}$ at low values is the result of plotting the distribution over $\log_{10}F_{\varphi}$.
@@
[…]
@@ IV.3 Joint constraints
[…]
@@ V.2 Local and late-time searches
In Section III we extend the analysis of these late-time probes to handle several qualitatively distinct features of hyperlight scalars, with results summarized in Fig. 2. Measurements long after a hyperlight scalar begins oscillating are plagued by its extremely long oscillation period: the statistical possibility that the measurement occurs near a node of the signal steeply penalizes the reach of single, finite-duration probes. As a result, atomic clocks and tests of the equivalence principle are limited in constraining power.888 The penalty does not apply to the discovery reach of such searches; see Section III.1 for further discussion. Section III.1 argues that the incurred penalty depends on the signal’s scaling with phase about its nodes. Equivalence principle tests (Section III.4) for quadratic couplings are therefore penalized by two powers of mass times inverse experimental duration, while atomic clock and pulsar timing arrays (Section III.6), which essentially probe the drift rate of fundamental constants, incur a penalty of one power for both linear and quadratic couplings. Combinations of measurements can mitigate this penalty only if spaced in time by a substantial fraction of the scalar’s oscillation period because of the greatly suppressed chance that the scalar’s amplitudes sampled at random points in its oscillation all occur at nodes. Constraints on varying constants from quasar absorption lines (Section III.3) span redshifts as high as 777 and therefore remain competitive when analyzed as an ensemble (Fig. 3), even if during epochs when the field oscillates, setting the leading constraints at the lowest masses.
@@ V.3 Cosmological signatures and constraints
Furthermore, this work provides theoretical grounding for the substantial body of literature constraining cosmological variations in fundamental constants on a phenomenological basis. The framework we develop, in the style of effective field theory, enables a relatively general comparison of the constraining power from cosmological data and independent laboratory and astrophysical searches. Prior work noted that linearly coupled scalars in principle do not redshift sufficiently after recombination to satisfy atomic clock constraints and simultaneously yield observably large variations in the CMB [191]. Our work shows that such arguments apply only to linearly coupled scalars that do not oscillate at late times: the fundamental-constant variation (or oscillation amplitude thereof) due to higher-order couplings redshifts faster, and marginalizing over the scalar’s unknown oscillation phase parametrically penalizes limits from terrestrial probes (see Section III.1). On the other hand, equivalence principle limits already entirely exclude any observable cosmological signatures from linearly coupled scalars (Section III.4), largely independent of the scalar’s abundance and potential [in the regime of interest with $V(\phi)\ll\mathrm{eV}^{4}$].
To assess the implications of laboratory and astrophysical searches for cosmological signatures, we may straightforwardly extrapolate the noncosmological limits presented in Fig. 5 to limits on variations at recombination. In the mass range for which the scalar remains frozen at recombination, the early-time variation in fundamental constants is $\sim d_{\lambda}^{(2)}F_{\varphi}$ [Eq. 58]. The limits reported in Fig. 5, taking fiducial $F_{\varphi}=10^{-2}$, thus correspond to the percent variation allowable at recombination. Limits from quasars preclude subpermille early-time variations only for $m_{\varphi}\lesssim 10^{-31}\leavevmode\nobreak\ \mathrm{eV}$ and quickly fall off at heavier masses. The limits from nucleosynthesis [Eq. 93] directly apply at recombination, so long as the scalar is static in the intervening epoch—i.e., if neither the bare nor in-medium potentials are important. The latter holds for scalar abundances near our fiducial choice, but for much less abundant scalars (at fixed early-time variation in fundamental constants) the dimensionless coupling, and therefore in-medium effects, are enhanced and nontrivially modify these conclusions. In summary, while this work establishes minimal models for which early-time variations in fundamental constants (including those motivated as early-recombination solutions to the Hubble tension) are viable in light of both theory and independent searches, these considerations essentially require it makes up a nonnegligible fraction of the total matter density today (the regime studied in Ref. [60]).
@@ V.4 Future directions
[…]
@@ A.1 After annihilation
Figure 6: Solutions to the Klein-Gordon equation Eq. 96 for a quadratically coupled scalar with a nonrelativistic matter potential, with values of $(Q_{b})_{\lambda}d_{\lambda}^{(2)}$ varying by color. Vertical black lines indicate the moment of recombination (i.e., for the fiducial case that the fundamental constants do not change with time). The right panel depicts the fractional difference in the shift in $\lambda$ from the initial condition, i.e., $|\lambda/\lambda_{i}-1|$.
@@
[…]
@@
where $F_{\varphi}$ is the scalar-CDM ratio given by Eq. 53, i.e., for an uncoupled scalar. Equation 102 encodes the correction to the relationship between the early-time parameter shift $\Delta\lambda_{i}/\lambda(0)$, $d_{\lambda}^{(2)}$, and the late-time abundance $F_{\varphi}$ for an uncoupled scalar, Eq. 58. As explained in Section IV.3, because these effects are relevant only at large $d_{\lambda}^{(2)}$ but cosmological data place relatively strong constraints on $\Delta\lambda_{i}/\lambda(0)$, they are only relevant in the regime in which the scalar’s abundance is negligibly small. Any changes to the scalar’s dynamics are then themselves negligible. Equation 102 still encodes the correction between $F_{\varphi}$ [taken to be defined as the combination in Eq. 58] and the present-day scalar abundance relevant to, e.g., laboratory probes. But these effects are again small in the parameter space we consider.
@@ A.2 Before annihilation
Figure 7: Solutions to the Klein-Gordon equation Eq. 103 for a scalar coupled quadratically to the electron (left) and photon (right), with couplings $d_{\lambda}^{(2)}$ varying by color.
@@
Oscillatory solutions for positive couplings again switch to exponentially growing ones for negative couplings. For large negative couplings $-d_{m_{e}}^{(2)}\gtrsim 100$, numerical solutions determine that the scalar refreezes at a value $\sim e^{3\sqrt{|\mathcal{D}_{m_{e}}|}}$ times larger than $\varphi_{\mathrm{RH}}$; the dependence on $\sqrt{|\mathcal{D}_{m_{e}}|}$ is as expected from Eq. 104, while the numerical coefficient reflects the particular dynamics of refreezing as the in-medium mass drops exponentially during annihilation. At the classical level, there is no theoretical issue with the scalar having an arbitrarily small value before reheating that is substantially amplified to its post-annihilation value. However, during inflation, a quantized scalar scalar field with $m_{\varphi}<H_{\text{inf}}$ acquires an average amplitude of order $\sim H_{\mathrm{inf}}^{2}/m_{\varphi}>H_{\text{inf}}$ [229]. Heuristically, this is because an inflationary Universe has a IR cutoff at the inflationary scale $H_{\text{inf}}$, meaning that the smallest physically meaningful scale is $\mathcal{O}(H_{\text{inf}})$. At the very minimum, inflation had to occur in time for reheating to complete at a temperature $T_{\mathrm{RH}}\approx 4\leavevmode\nobreak\ \mathrm{MeV}$ [85, 86, 87, 88, 89, 90, 91], limiting $H_{\mathrm{inf}}\gtrsim 10^{-14}\leavevmode\nobreak\ \mathrm{eV}$. Requiring that $|\varphi_{\text{RH}}|>H_{\text{inf}}/\sqrt{2}M_{\mathrm{pl}}$ implies
@@
The scalar’s effective mass induced by its coupling to photons is proportional to $H$ at all times after reheating.101010 Technically, the rest of the charged SM content would contribute to the interacting free energy at higher temperatures. The SM presumably also generates an effective potential for the scalar before it fully thermalizes, but it would not be captured by the thermal field theory calculations Section II.3.2 relies on. We set aside these complex and uncertain details for the sake of simplicity. Now defining $\mathcal{D}_{e}=25\alpha(0)d_{e}^{(2)}/8\pi g_{\star}$, the solution under the photon coupling is
@@
where $\Delta\alpha_{i}$ is the change in $\alpha$ post-annihilation that is the initial condition for the dynamics leading up to recombination. Again taking the minimum reheat temperature $T_{\mathrm{RH}}\approx 4\leavevmode\nobreak\ \mathrm{MeV}$, this bound amounts to $d_{e}^{(2)}\lesssim 360$; for instantaneous reheating after high-scale inflation, the bound tightens by a factor 20$\sim$ 20∼ 20.
@@ B.1 Impact of priors on marginalized constraints
[…]
@@
Compared to a uniform prior over $F_{\varphi}$ and $\lambda_{i}/\lambda_{0}$, this prior penalizes large $F_{\varphi}$ quadratically. If the posteriors over $d_{\lambda}^{(2)}$ are largely driven by the likelihood’s dependence on $\lambda_{i}/\lambda_{0}$, we would expect such a prior to drive the marginal posteriors over the quadratic couplings $d_{\lambda}^{(2)}$ to values even larger in magnitude. Priors must be proper, and the chosen limits on, e.g., $\bar{\varphi}_{i}$ determine the relative prior weight allocated to percent-level values of $F_{\varphi}$—i.e., those that were allowed in the results taking uniform phenomenological priors. Specifically, we impose the effective prior Eq. 113 with boundaries corresponding to $d_{\lambda}^{(2)}\sim\mathcal{U}(-10^{3},10^{3})$ and $\log_{10}\bar{\varphi}_{i}\sim\mathcal{U}(-2,0)$, where $\mathcal{U}$ denotes a uniform distribution. We additionally restrict $10^{-4}\leq F_{\varphi}\leq 0.3$ in order to roughly match the range of values covered by the posteriors using uniform, phenomenological priors.